# An $\tilde{\mbox{O}}$ptimal Differentially Private PAC Learner for Concept Classes with VC Dimension 1

### v1

@@ Abstract
We present the first nearly optimal differentially private PAC learner for any concept class with VC dimension 1 and Littlestone dimension $d$. Our algorithm achieves the sample complexity of $\tilde{O}_{\varepsilon,\delta,\alpha,\delta}(\log^{*}d)$, nearly matching the lower bound of $\Omega(\log^{*}d)$ proved by Alon et al. [1] [STOC19]. Prior to our work, the best known upper bound is $\tilde{O}(VC\cdot d^{5})$ for general VC classes, as shown by Ghazi et al. [13] [STOC21].
@@ 1 Introduction
Differentially private learning, introduced by Kasiviswanathan et al. [17], studies the task of learning a hypothesis from data while preserving the privacy of individual entries in the dataset. Formally, the goal is to construct a learner that satisfies the requirements of PAC learning and, simultaneously, differential privacy.
Definition 1 (PAC learning [21]).
Given a set of $n$ points $S=(\mathcal{X}\times\{0,1\})^{n}$ drawn i.i.d. from an unknown distribution $\mathcal{D}$ and labels that are given by an unknown concept $c\in\mathcal{C}$, we say a learner $L$ (possibly randomized) $(\alpha,\beta)$-PAC learns $\mathcal{C}$ if $h=L(S)$ and
$\Pr[error_{\mathcal{D}}(c,h)\leq\alpha]\geq 1-\beta.$
@@
Definition 3 (differentially private learning [17]).
We say a learner $L$ $(\alpha,\beta,\varepsilon,\delta)$-differentially privately PAC learns the concept class $\mathcal{C}$ if
  1. 1.
    $L$ $(\alpha,\beta)$-PAC learns $\mathcal{C}$.
  2. 2.
    $L$ is $(\varepsilon,\delta)$-differentially private.
We call the dataset size of the learning task the sample complexity, which is a fundamental question in learning theory. For non-private PAC learning, it is well-known that the sample complexity is linear to the VC dimension of the concept class [21]. However, the sample complexity is less well understood in the differentially private setting. Kasiviswanathan et al. [17] define the private learning and give a general upper bound $O(\log|\mathcal{C}|)$, which works for pure differential privacy ($\delta=0$). This bound is tight for several natural concept classes [2, 12].
In the approximate privacy regime ($\delta>0$), several works [4, 8, 3, 16, 15, 9, 19] show that the sample complexity can be significantly lower than that in the pure setting. Alon et al. [1] and Bun et al. [7] find that the sample complexity of approximately differentially private learning is related to the Littlestone dimension [18] of the concept class $\mathcal{C}$. In detail, for a concept class $\mathcal{C}$ with Littlestone dimension $d$, Alon et al. [1] prove a lower bound $\Omega(\log^{*}d)$ and Bun et al. [7] provide an upper bound $\tilde{O}(2^{2^{d}})$. Subsequently, Ghazi et al. [13] improve the upper bound to $\tilde{O}(VC\cdot d^{5})$.
This leaves a large gap between the lower bound $\Omega(VC+\log^{*}d)$ and the upper bound $\tilde{O}(VC\cdot d^{5})$, even when the VC dimension is as small as 1. The main question is: what is the correct dependence on $d$? One important example is the halfspaces class, which can be privately learned with $poly(VC,\log^{*}d)$ examples by the work of Nissim et al. [19]. So it’s natural to ask: {mdframed} Could we privately learn any concept class with sample size $poly(VC,\log^{*}d)$?
Unfortunately, the technique of Nissim et al. [19] cannot extends to the general VC class because it depends on the data structure of halfspaces. However, the work by Nissim et al. [19] also shows an upper bound of $\tilde{O}(\log^{*}|\mathcal{X}|)$ for any VC 1 class. Although $|\mathcal{X}|$ can be infinitely larger than $d$ for the general concept class, and a significant gap remained between $\log^{*}d$ and $\min\{d^{5},\log^{*}|\mathcal{X}|\}$, it shows that in some cases (say when $\log|\mathcal{X}|=d$), the sample complexity of of private learning depends on $\log^{*}d$.
To the target of fully understand the sample complexity of differentially private learning, the first step is to ask:
{mdframed}
Could we privately learn VC 1 class with sample size $poly(\log^{*}d)$?
@@ 1.2 Overview of Technique
The key observation is that any concept with VC dimension 1 has a tree structure, which is observed by Ben-David [5]. In the tree structure, each node is a point from the domain $\mathcal{X}$, and each hypothesis is a path from one node to the tree’s root. We show that the tree’s height is upper bounded by the threshold dimension of the concept class, which has an upper bound $O(2^{d})$ [20, 14, 1].
We use the partition and aggregate method to construct the private learner. The labeled dataset $S$ is randomly partitioned into subsets $S_{1},\dots,S_{t}$, where $t=O(\log^{*}d)$. Each subset has a set of ”deterministic points”, whose labels are fixed to be 1 by the given labeled points. We show that the deterministic points can be used to construct an accurate hypothesis, and all the sets of deterministic points of $S_{1},\dots,S_{t}$ are on the same path. Since the length of the path is at most $O(2^{d})$, we use the private median algorithm [9] to select a ”good length” with sample size $\tilde{O}(\log^{*}d)$. Then we can use the choosing mechanism [4] to select a ”good path” with the ”good length” with a sample size $O(1)$ because the deterministic points of $S_{1},\dots,S_{t}$ are on the same path. Finally, we show that the selected ”good path” is an accurate hypothesis.
@@ 3.1 Learning theory
For a domain $\mathcal{X}$ and a concept class $\mathcal{C}$, We say $x_{1},\dots,x_{k}$ is shattered if for any subset $S\subseteq\{x_{1},\dots,x_{k}\}$, there exists a concept $c\in\mathcal{C}$, such that $c(x)=1$ for any $x\in S$ and $c(x)=0$ for any $x\in\{x_{1},\dots,x_{k}\}\backslash S$. The maximum size of the shattered set is called VC dimension.
@@ 3.2 Differential privacy
For any set of number $S=\{x_{1},\dots,x_{n}\}$, there exists a $1/2$-median.
@@ 4 Structure of Classes with VC Dimension 1
Given the pair of the point and label $(x,c(x))$ that is labeled by $c$, we can transform it to a corresponding pair labeled by $c_{f}$: let label be 1 if $c(x)\neq f(x)$ and be 00 if $c(x)=f(x)$.
@@
      output $1$ if $y\neq f(x_{new})$, otherwise output 0.
@@ 4.1 Tree Structure
5 Private learner
Parameter: Confidence parameter $\beta>0$, privacy parameter $\varepsilon,\delta>0$, $n_{PM}(d,1/3,\beta,\varepsilon,\delta)=O\left(\frac{\log^{*}d\cdot\log^{2}(\frac{\log^{*}d}{\beta\delta})}{\varepsilon}\right)$ and number of subsets $t=\max\left\{n_{PM}(d,1/3,\beta,\varepsilon,\delta),O\left(\frac{1}{\varepsilon}\log(\frac{4n_{PM}(d,1/3,\beta,\varepsilon,\delta)}{\beta\varepsilon\delta})\right)\right\}=O\left(\frac{\log^{*}d\cdot\log^{2}(\frac{\log^{*}d}{\varepsilon\beta\delta})}{\varepsilon}\right)$, where $d=TD({\cal X},\mathcal{C}_{f})+1$.
@@ 5 Private learner
    construct tree according to Algorithm 1
@@
    compute the $1/3$-median $z=PrivateMedian(y_{1},\dots,y_{t})$ with parameter $\varepsilon,\delta,\beta$.
@@
Notice that $B_{i}\subseteq I(h^{\prime})$. Consider the set $I(h^{\prime})\backslash B_{i}$. For all $x\in(h_{1}\backslash B_{i})$, we have $c^{*}_{f}(x)=0$ because $x\notin I(c^{*}_{f})$. Therefore $error_{\mathcal{D}}(h_{i},c^{*}_{f})\leq error_{\mathcal{D}}(h^{\prime},c^{*}_{f})\leq\alpha$ (because for all points that $h^{\prime}$ and $h_{i}$ make different predictions, $h_{i}$ gives the correct label).

### v2

@@ Abstract
We present the first nearly optimal differentially private PAC learner for any concept class with VC dimension 1 and Littlestone dimension $d$. Our algorithm achieves the sample complexity of $\tilde{O}_{\varepsilon,\delta,\alpha,\beta}(\log^{*}d)$, nearly matching the lower bound of $\Omega(\log^{*}d)$ proved by Alon et al. [1] [STOC19]. Prior to our work, the best known upper bound is $\tilde{O}(VC\cdot d^{5})$ for general VC classes, as shown by Ghazi et al. [13] [STOC21].
@@ 1 Introduction
Machine learning algorithms can access sensitive information from the training dataset. We research the privacy-preserving machine learning technique, introduced by Kasiviswanathan et al. [17], that targets to learn a hypothesis while preserving the privacy of individual entries in the dataset. Informally, the goal is to construct a learner that satisfies the requirements of probably approximately correct (PAC) learning [22] and, simultaneously, differential privacy [11].
@@
In the PAC learning task, the learner is given a set of $n$ points $S=(\mathcal{X}\times\{0,1\})^{n}$ drawn i.i.d. from an unknown distribution $\mathcal{D}$ and labels that are given by an unknown concept $c:\mathcal{X}\rightarrow\{0,1\}$, where $c\in\mathcal{C}$ and $\mathcal{C}$ is the concept class. The goal is to output a hypothesis $h:\mathcal{X}\rightarrow\{0,1\}$ such that $\Pr_{x\sim\mathcal{D}}[c(x)\neq h(x)]\leq\alpha$, where $\alpha$ is the parameter of accuray. We call the required dataset size of the learning task the sample complexity, which is a fundamental question in learning theory. For non-private PAC learning, it is well-known that the sample complexity is linear to the VC dimension of the concept class [22]. However, the sample complexity is less well understood in the differentially private setting. Kasiviswanathan et al. [17] give a upper bound $O(\log|\mathcal{C}|)$, which works for pure differential privacy ($\delta=0$).
[…]
This leaves a large gap between the lower bound $\Omega(\log^{*}d)$ and the upper bound $\tilde{O}(VC\cdot d^{5})$, even when the VC dimension is as small as 1. The main question is: what is the best dependence on $d$? One important example is the halfspaces class, which can be privately learned with $poly(VC,\log^{*}d)$ examples by the work of Nissim et al. [20]. So it’s natural to conjecture that the tight sample complexity of differentially private learning depends on $\log^{*}d$.
Moreover, the work by Nissim et al. [20] also shows an upper bound of $\tilde{O}(\log^{*}|\mathcal{X}|)$ (recall $|\mathcal{X}|$ is the input domain size) for any concept class with VC dimension 1. However, $|\mathcal{X}|$ could be arbitrarily larger than $d$ for the general concept class, and a significant gap remained between $\log^{*}d$ and $\min\{d^{5},\log^{*}|\mathcal{X}|\}$. The remaining question is:
Could we privately learn any VC 1 class with sample size $poly(\log^{*}d)$?
@@ 1.1 Our Result
Theorem 2 (Proper Learning).
$\begin{array}[]{rl}N&\geq O\left(\frac{\log^{*}d\cdot\log^{2}(\frac{\log^{*}d}{\varepsilon\beta\delta})}{\varepsilon}\cdot\frac{48}{\alpha}\left(10\log(\frac{48e}{\alpha})+\log(\frac{5}{\beta})\right)+\frac{(\log(1/\alpha)+\log(1/\alpha\beta))\cdot\sqrt{\log(1/\delta)}}{\alpha^{2.5}\varepsilon}\right)\\ &=\tilde{O}_{\beta,\delta}\left(\frac{\log^{*}d}{\alpha\varepsilon}+\frac{1}{\alpha^{2.5}\varepsilon}\right)\end{array}$
there is an $(\varepsilon,\delta)$-differentially private algorithm that properly $(\alpha,\beta)$-PAC learns $\mathcal{C}$.
1.2 Other Related Works
Two Well Researched VC 1 Classes
The first class is the point functions class (also called the indicator function). It is defined to be $f_{t}(x)=1$ if $x=t$ and $f_{t}(x)=0$ if $x\neq t$. The point functions class has Littlestone Dimension 1, and Beimel et al. [4] prove it can be privately learned with $O(1)$ examples using the choosing mechanism (Lemma 1).
The second example is the thresholds class. It is defined to be $f_{t}(x)=1$ if $x\geq t$ and $f_{t}(x)=0$ if $x<t$. The thresholds class has Littlestone dimension $\log|\mathcal{X}|$. There is a series of works on privately learning thresholds [4, 8, 15, 9], and finally been found that we can privately learn thresholds with $\tilde{O}(\log^{*}|\mathcal{X}|)$ examples. The key observation is the equivalence between private learning thresholds and private median (or private interior point), which is observed by Bun et al. [8].
Proper and Improper Private Learning
We call a learner proper if the output hypothesis is in the concept class and improper if the output may not be in the concept class. For pure differential privacy, Beimel et al. [2] prove a lower bound on the sample complexity of proper learning point functions is $\Omega(\log|X|)$ and an upper bound of improper learning point functions $O(1)$. It shows a significant separation between proper and improper learning with pure differential privacy. Our result shows that for any VC 1 class, the separation between proper and improper learning with approximate differential privacy cannot be larger than $O(\frac{1}{\alpha^{O(1)}})$ (ignoring other parameters).
@@ 1.3 Overview of Technique
1.3.1 Improper Learning
The key observation is the tree structure [^2] The Tree structure of the VC 1 class is mentioned in a lot of papers and is usually researched in the one-inclusion graph context. But most of them only consider the maximum concept classes. The observation by Ben-David can include non-maximum classes. and partial order in the VC 1 class, which is observed by Ben-David [5]. In the tree structure, each node is a point from the domain $\mathcal{X}$, and each hypothesis is a path from one node to the tree’s root. Another important idea is that we focus on the thresholds dimension instead of the Littlestone dimension. We show that the tree’s height is upper bounded by the threshold dimension of the concept class, which has an upper bound $O(2^{d})$ [21, 14, 1]. So that we can apply the private median algorithm (Theorem 9) and achieve $\tilde{O}(\log^{*}d)$ sample complexity.
Based on the tree structure, we show that every VC 1 class can be privately learned in two steps:
Step 1: privately find a ”good length”.
We use the partition and aggregate method to construct the private learner. The labeled dataset $S$ is randomly partitioned into subsets $S_{1},\dots,S_{t}$, where $t=\tilde{O}(\log^{*}d)$. Each subset has a set of ”deterministic points”, whose labels are fixed to be 1 by the given labeled points. We show that the deterministic points can be used to construct an accurate hypothesis, and all the sets of deterministic points of $S_{1},\dots,S_{t}$ are on the same path. Since the length of the path is at most $O(2^{d})$, we use the private median algorithm (Theorem 9 [9]) to select a ”good length” with sample size $\tilde{O}(\log^{*}d)$.
Step 2: privately find a ”good path”.
We let the ”good path” be the path such that all points on this path are deterministic by the given subset. We can show that for all paths with the ”good length”, only one can be the required ”good path”. So that we transform it to be a 1-bounded function (Definition 12) and we can use the choosing mechanism (Lemma 1) to select a ”good path”. This ”good path” can be projected to an accurate hypothesis.
@@ 1.3.2 Example
1.3.3 Proper Learning
The key observation is that the improper learner (Algorithm 2) outputs $x_{good}$, which is on the path of the true concept path. Let $x_{c^{*}_{f}}$ be the corresponding node of the true concept. Then, for all nodes on the path from $x_{good}$ to $x_{c^{*}_{f}}$, they can project to a highly accurate hypothesis. We call a node a proper node if it can be projected to a concept in $\mathcal{C}$. Thus, we can only consider the sub-tree $T_{good}$ that satisfies (see example in Figure 3):
  1. 1.
    the root of $T_{good}$ is $x_{good}$
  2. 2.
    all leaves of $T_{good}$ are proper nodes
  3. 3.
    all nodes except leaves are not proper nodes
Figure 3: We use the circle to represent the improper node and the star to represent the proper node. Let the red node be $x_{good}$. The figure shows the subtree with root $x_{good}$.
By the above observation, there must exist one leaf on the path from $x_{good}$ to $x_{c^{*}_{f}}$, and thus it has high accuracy and can be projected to a concept in $\mathcal{C}$. By the analysis of the improper learner, for all nodes on the sub-tree with root $x_{good}$, they correctly classify most points with label 1. Thus, we only need to consider the accuracy on the points with label 0. If we can find a leaf $x^{\prime}$ such that only a small number of points with label 0 are located on the path from $x_{good}$ to $x^{\prime}$, then for the corresponding hypothesis $h^{\prime}$, where $I(h^{\prime})=\{x|x^{\prime}\preceq x\}$, $h^{\prime}$ and $h_{good}$ (recall that $h_{good}$ is corresponding hypothesis of $x_{good}$) disagree a small number of labeled points in $S$. Thus, $h^{\prime}$ has high accuracy.
Let $v(x)$ be the number of points with label 0 on the path from $x$ to $x_{good}$. The target is to find a proper node $x$ with a small value $v(x)$. Let $w(x)$ be the number of points with label 0 below $x$ on $T_{good}$. Notice that for all $x^{\prime}\prec x$, we have $v(x^{\prime})\leq v(x)+\sum_{x^{\prime}\in T_{x}}[w(x)]$, where $T_{x}$ is sub-tree in $T_{good}$ with root $x$. So that we can look for a point $x$ such that $v(x)+\sum_{x^{\prime}\in T_{x}}[w(x^{\prime})]$ is small. Then, all leaves of $T_{x}$ are proper nodes with high accuracy. Consider another dataset with size $N$ and the following strategy to look for it:
  1. 1.
    we set $x_{flag}=x_{good}$ at the beginning, let $X_{children}$ be the set of all children of $x_{flag}$.
  2. 2.
    Case 1: there exists $x\in X_{children}$ such that $w(x)\leq O(\alpha\cdot N)$, then all leaves $\ell$ of sub-tree $T_{x}$ have $v(\ell)\leq w(x)+v(x)$. In this case, we cannot have too many ”bad children” with heavy weights because every pair of children does not have an intersection part on their subtree, and the sum of weights cannot be larger than the number of points. Running the exponential mechanism can privately select a child with light weight even though $|X_{children}|$ may be large.
  3. 3.
    Case 2: for any $x\in X_{children}$, we have $w(x)>\Omega(\alpha\cdot N)$. Since $\sum_{x\in X_{children}}w(x)\leq N$, we have $|X_{children}|\leq O(\frac{1}{\alpha})$. So we can use the exponential mechanism to select a child, such that it contains a ”good leaf” $\ell$ satisfying $v(\ell)$ is small. Since $|X_{children}|\leq O(\frac{1}{\alpha})$, the error is bounded by $O(\frac{\log(1/\alpha)}{\varepsilon})$, where $\varepsilon$ is the privacy parameter.
In case 1, we immediately find the good leaf. In case 2, we remove at least $\alpha\cdot N$ points that are located in the subtree of the other children. Thus, we can stop in $O(\frac{1}{\alpha})$ steps and the accumulated error is bounded by $O(\frac{\log(1/\alpha)}{\alpha\varepsilon})$. That is, in each iteration, there must exist a leaf $\ell$ in the subtree $T_{x_{flag}}$ satisfying $v(\ell)\leq\min_{\ell^{\prime}\in T_{good}}v(\ell^{\prime})+O(\frac{\log(1/\alpha)}{\alpha\varepsilon})$. Since running case 1 causes error $O(\alpha\cdot N)$ and the improper learner guarantees that $\min_{\ell^{\prime}\in T_{good}}v(\ell^{\prime})\leq O(\alpha\cdot N)$, the above strategy will make $v(\ell)\leq O(\alpha\cdot N+\frac{\log(1/\alpha)}{\alpha\varepsilon})$.
@@ 3.1 Learning theory
Definition 4 (Proper Learning).
For the concept class $\mathcal{C}$, we call the learner proper if the output hypothesis $h\in\mathcal{C}$. We call the learner improper if the output hypothesis $h\notin\mathcal{C}$.
@@
[…]
@@ 3.2 Differential privacy
Theorem 5 (Post-processing [11]).
Theorem 6 (Composition [10]).
Theorem 7 (Advanced composition [12]).
Let $M_{1},\dots,M_{k}:{\cal X}^{n}\rightarrow{\cal Y}$ be $(\varepsilon,\delta)$-differentially private mechanisms. Then the algorithm that on input $S\in{\cal X}^{n}$ outputs $(M_{1}(S),\dots,M_{k}(S))$ is $(\varepsilon^{\prime},k\delta+\delta^{\prime})$-differentially private, where $\varepsilon^{\prime}=\sqrt{2k\ln(1/\delta^{\prime})}\cdot\varepsilon$ for every $\delta^{\prime}>0$.
Theorem 8 (Exponential Mechanism [19]).
[…]
@@
For any set of number $S=\{x_{1},\dots,x_{n}\}$, there exists a 1/21/21 / 2-median.
@@ 4 Structure of Classes with VC Dimension 1
Given the pair of the point and label $(x,c(x))$ that is labeled by $c$, we can transform it to a corresponding pair labeled by $c_{f}$: let label be 1 if $c(x)\neq f(x)$ and be 0 if $c(x)=f(x)$.
@@
      output 111 if $y\neq f(x_{new})$, otherwise output 0.
@@ 4.1 Tree Structure
5 Improper learner
Parameter: Confidence parameter $\beta>0$, accuracy paramenter $\alpha>0$, privacy parameter $\varepsilon,\delta>0$, $n_{PM}(d,1/3,\beta,\varepsilon,\delta)=O\left(\frac{\log^{*}d\cdot\log^{2}(\frac{\log^{*}d}{\beta\delta})}{\varepsilon}\right)$ and number of subsets $t=\max\left\{n_{PM}(d,1/3,\beta,\varepsilon,\delta),O\left(\frac{1}{\varepsilon}\log(\frac{4n_{PM}(d,1/3,\beta,\varepsilon,\delta)}{\beta\varepsilon\delta})\right)\right\}=O\left(\frac{\log^{*}d\cdot\log^{2}(\frac{\log^{*}d}{\varepsilon\beta\delta})}{\varepsilon}\right)$, where $d=TD({\cal X},\mathcal{C}_{f})+1$.
@@ 5 Improper learner
    construct tree according to Algorithm 1, let $f$ be the select concept in Algorithm 1, transfer all labels in $S$ according to $f$ (by Definition 14).
@@
    compute the 1/31/31 / 3-median $z=PrivateMedian(y_{1},\dots,y_{t})$ with parameter $\varepsilon,\delta,\beta$.
@@
[…]